OpenStudy (anonymous):

Find the coordinates of the center of the following circle. (x - 5)² + (y + 3)² = 25 (-5, 3) (5, -3)

OpenStudy (xapproachesinfinity):

if you have \((x-a)^2+(x-b)^2=r^2\) the center is (a,b) and the radius is r

OpenStudy (xapproachesinfinity):

look at you circle equation and see what you need to do

OpenStudy (xapproachesinfinity):

my mistake it should be with y-b not x-b

OpenStudy (xapproachesinfinity):

\((x-a)^2+(y-b)^2=r^2\) like this now it looks like your equation look for a and b

OpenStudy (xapproachesinfinity):

do you get it

OpenStudy (anonymous):

not really

OpenStudy (xapproachesinfinity):

every circle equation is written like \((x-a)^2+(y-b)^2=r^2\) this is a general circle equation with center (a,b) in your case you have \((x-\color{red}{5})^2+(y\color{red}{+}\color{blue}{3})^2=25\)

OpenStudy (xapproachesinfinity):

a=5 and b=-3 in your case

OpenStudy (xapproachesinfinity):

since we have y+3 that means y-(-3) that's b=-3 so the center is (5,-3)

OpenStudy (anonymous):

ohh got it thnx

OpenStudy (xapproachesinfinity):

welcome! try finiding the center of this circle \(x^2+(y-2)^2=36\)